OpenStudy (anonymous):

What is the period, in seconds, of a simple pendulum of length 5 meters? Use the gravitational constant g = 9.8 m/s2 and round your answer to two decimal places.

OpenStudy (anonymous):

\[(\frac1{2\pi})\sqrt{\frac {l}{g}}\]

OpenStudy (anonymous):

.11?